A can solve 90 percent of the problems given in a book and B can solve 709 percent. what is the probability that at least one of them will solve a problem selected at random?
1
Expert's answer
2020-10-27T17:07:24-0400

Let A and B be the events defined as follows :

"A" be the event "A solves the problem", "B" be the event "B solves the problem",

 "A"  and "B"  are independent events such that


"P(A)=0.9, P(B)=0.7"

The probability that at least one of them will solve a problem selected at random


"P(A\cup B)=1-P(\bar{A})P(\bar{B})="

"=1-(1-P(A))(1-P(B))="

"=1-(1-0.9)(1-0.7)=0.97"

Or


"P(A\cup B)=P(A)+P(B)-P(A\cap B)="

"=P(A)+P(B)-P(A)P(B)="

"=0.9+0.7-0.9(0.7)=0.97"